OpenStudy (anonymous):

equation in interval [0,2pi) 10. sin 2x + cos x = 0

OpenStudy (jamesj):

Well, sin 2x = 2 sin x . cos x See what to do now?

OpenStudy (jamesj):

No? Well this means that sin 2x + cos x = 0 is equivalent to 2.sin x.cos x + cos x = 0 cos x (2sin x + 1) = 0 Now see what to do?

OpenStudy (anonymous):

The problem looks for pi something

OpenStudy (anonymous):

Intervals such as 0,2pi

OpenStudy (jamesj):

If ab = 0, what can you say about a and b? If ab = 0, then a = 0 or b = 0 Likewise if cos x (2sin x + 1) = 0 then cos x = 0 or 2sin x + 1 = 0 Now solve for x, given that x is in the interval [0, 2pi)

OpenStudy (anonymous):

I don't how

OpenStudy (jamesj):

Well if cos x = 0, what must x be if x is in [0, 2pi) ?

OpenStudy (anonymous):

I don't understand